OpenStudy (anonymous):

Which expression is equivalent to (8^2)−4?

undeadknight26 (undeadknight26):

May I have some expressions?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

anyone there ?

OpenStudy (anonymous):

*stretches panda arms and yawns*

undeadknight26 (undeadknight26):

LOL The role play. I'm checking it out: Right now I used a calculator to find out that (8^2)^-4 = 1/16777216

OpenStudy (anonymous):

omg ive never been so confuzed. my panda brain is small.

undeadknight26 (undeadknight26):

LOL ISH OKAI KAWAII CREATURE! I SHALL SOLVE THIS!

OpenStudy (anonymous):

ᕦ( ͡° ͜ʖ ͡°)ᕤ

OpenStudy (anonymous):

lmao

undeadknight26 (undeadknight26):

@iambatman ;~; Plz help ;~: I confuzzled too ;~;

OpenStudy (anonymous):

@llamb25

OpenStudy (anonymous):

\[(x^m)^n \implies x^{mn}\] and you will need \[x^{-n} = \frac{ 1 }{ x^n }\] So we have \[\large (8^2)^{-4} = 8^{2 \times -4} = 8^{-8}\] now use the property above and remember \[8^8 = 8 \times 8 \times 8 \times 8 \times 8 \times 8 \times 8 \times 8\]

OpenStudy (anonymous):

ok so 1 ---------------------------------- 88888888

OpenStudy (anonymous):

yay for pandas!

OpenStudy (anonymous):

so its C?

OpenStudy (anonymous):

Sounds good!

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

thanks